OpenStudy (samsan9):

give an example of how the zero product property applies to solving quadraticequations

OpenStudy (precal):

(x+a)(x+b)=0 x=-a x=-b

OpenStudy (precal):

I believe the zero product property allows you to set each factor to zero and solve it. Hence the name zero product property. Product means times example (x+3)(x-2)=0 x+3=0 x-2=0 x=-3 x=+2